OpenStudy (logan13):

If the diameter of the circle above is 20 mm, what is the area of the circle? 100 mm2 20 mm2 400 mm2 10 mm2

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (fibonaccichick666):

So, do you know the formula for area of a circle?

OpenStudy (fibonaccichick666):

That would be a picture of the diameter, not the formula for area

OpenStudy (alexandervonhumboldt2):

radius=1/2 diameter are=radius^2*3.14

OpenStudy (logan13):

c

OpenStudy (unanimoose):

Ok so like I said before your options probably have the pi symbol in them so you don't actually have to multiply by 3.14 here.

OpenStudy (unanimoose):

Jsut divide the diameter by 2 and square it.

OpenStudy (unanimoose):

Because: \[A = r^2\pi\]

OpenStudy (unanimoose):

Nope, it isn't C.